OpenStudy (anonymous):

can you tell me ifthis is right??? -32/85 / 50/51 -32/85 * 50/51 -32/17 * 10/51 -320/17*51 -320/867 -32/85 / 50/51= -320/867

OpenStudy (anonymous):

are those 2 fraction ?

OpenStudy (anonymous):

-32/85 / 50/51

OpenStudy (anonymous):

you make that -32/85 * 50/51

OpenStudy (anonymous):

ok ill take that as a yes, whenever you're dividing by a fraction, you can multiply by the inverse

OpenStudy (anonymous):

so is that right or no

OpenStudy (anonymous):

no basically 1/ (1/2) = 1*(2/1)

OpenStudy (anonymous):

ok i'm lost

OpenStudy (anonymous):

sorry i mean multiply by the recipocal not the inverse

OpenStudy (anonymous):

another example (3/4) / (5/8) = (3/4)*(8/5) this is because you multiply (8/5)/(8/5) ,which is also equal to 1, to the entire equation